# Boundary transfer matrices arising from quantum symmetric pairs

### v1

@@ 5. Quantum symmetric pairs and K-matrices
We provide a brief summary of the theory of quantum symmetric pairs and their K-matrices, and survey some recent works on related topics. The main references for proofs and further details in this section are [Kol14, RV22, AV22a, AV24].
In [AV22a, AV24], we constructed a family of (basic and tensor) K-matrices arising from quantum symmetric pairs, whose action is defined on ${{\mathcal{O}}^{\sf int}}$ and, in fact, on modules in ${\mathcal{O}}$ satisfying a weaker integrability condition. Henceforth, for simplicity of the exposition, we consider only the category ${{\mathcal{O}}^{\sf int}}$, and we set
@@ 5.1. Pseudo-involutions and their q-deformations
A Lie algebra automorphism $\theta:{\mathfrak{g}}\to{\mathfrak{g}}$ is said to be of the second kind if $\theta({\mathfrak{b}}^{+})$ is obtained by the adjoint action of the Kac-Moody group on the opposite Borel subalgebra ${\mathfrak{b}}^{-}$; this is equivalent to $\dim({\mathfrak{n}}^{+}\cap\theta({\mathfrak{n}}^{+}))<\infty$, see e.g., [KW92, 4.6]. If $\theta$ is an involution of the second kind then, up to conjugation by an inner automorphism, we may assume $\theta({\mathfrak{h}})={\mathfrak{h}}$ and $\theta$ fixes pointwise any $\theta$-stable root space, see [KW92, Ch. 5]. The main example is the Chevalley involution $\omega:{\mathfrak{g}}\to{\mathfrak{g}}$, defined by
@@
where $\tau$ is a diagram automorphism [^9] Recall that a diagram automorphism (a bijection $\tau:I\to I$ such that $a_{\tau(i)\tau(j)}=a_{ij}$ for all $i,j\in I$) acts as a Lie algebra automorphism of ${\mathfrak{g}}^{\prime}$ by relabelling: $\tau(e_{i})=e_{\tau(i)}$, $\tau(f_{i})=f_{\tau(i)}$. By [KW92, 4.19], there exists a complementary subspace ${\mathfrak{h}}^{\prime\prime}$ of ${\mathfrak{h}}^{\prime}={\mathfrak{h}}\cap{\mathfrak{g}}^{\prime}$ such that the action of $\tau$ can be extended to ${\mathfrak{g}}={\mathfrak{g}}^{\prime}\oplus{\mathfrak{h}}^{\prime\prime}$., $\operatorname{Ad}(\widetilde{w}_{X})$ is the triple exponential action induced by the longest element $w_{X}$ of a unique subdiagram $X\subseteq I$ of finite type and a character $\chi:\mathsf{Q}\to{\mathbb{C}}^{\times}$ such that $\chi(\alpha_{i})=1$ if $i\in X$. Note that $(X,\tau,\chi)$ are uniquely determined by $\phi$ from the expression (5.3). In particular, $X=\{i\in I\,|\,\phi(\alpha_{i})=\alpha_{i}\}$ (see [RV22, Sec. 2] and [AV22a, Sec. 6.1 & 6.2] for details).
In order for $\phi$ to lift to an algebra automorphism $\phi_{q}:U_{q}({\mathfrak{g}})\to U_{q}({\mathfrak{g}})$, it suffices that the extension of $\tau$ to ${\mathfrak{h}}^{\prime\prime}$ described by [KW92, 4.19] permutes the ${\mathbb{Z}}$-basis of $\mathsf{Q}^{\vee}_{\sf ext}$, constraining the choice of $d_{s}$ ($1\leq s\leq{\operatorname{cork}}(A)$). In the affine case, a suitable choice of ${\mathbb{Z}}$-basis can always be made, see [Kol14, Prop. 2.12], although not in general if ${\operatorname{cork}}(A)=1$, see [AV22a, Sec. 3.15].
In [AV22a, Sec. 6.7] a definition for the lift $\phi_{q}:U_{q}({\mathfrak{g}})\to U_{q}({\mathfrak{g}})$ is given, which can be related to the original constructions in [Let99, Kol14]. Note that here we include into $\phi_{q}$ the adjoint action of a lift of $\chi$ to a multiplicative character $\chi_{q}:\mathsf{Q}\to{\mathbb{F}}^{\times}$ such that $\chi_{q}(\alpha_{i})=1$ if $i\in X$. In terms of the formalism of unrestricted specialization, see e.g., [DCK90], the automorphism $\phi_{q}$ specializes to the automorphism $\phi$.
@@ 5.2. Pseudo-fixed-point subalgebras and their q-deformations
It is useful to constrain $\phi$ by imposing additional conditions on $X(\phi)$, $\tau(\phi)$, $\chi(\phi)$, leading to the notion of an (enriched) generalized Satake diagram, see [RV22, Def. 3.4] and [AV22a, Sec. 6.4]. The resulting conditions on $\phi$ are equivalent to requiring that the associated pseudo-fixed-point subalgebra
@@
It is convenient to allow some extra degrees of freedom and define a modification $U({\mathfrak{k}})_{s}$ of $U({\mathfrak{k}})$ by adding a scalar $s_{i}\in{\mathbb{C}}$ to the generator $b_{i}\in U({\mathfrak{k}})$ for $i\in I\backslash X$. In order to define the Letzter-Kolb deformation of $U({\mathfrak{k}})_{s}$, we choose a pseudo-involution $\phi$ with a canonical expression (5.3) in terms of an associated enriched generalized Satake diagram $(X,\tau,\chi)$. The induced constraint on $\chi_{q}$ is conveniently rewritten in terms the tuple $\gamma=(\chi_{q}(\alpha_{\tau(i)}))_{i\in I}$. Also, to deduce basic properties of the deformed algebra, we will need a constrained tuple $\sigma\in{\mathbb{F}}^{I}$ which specializes to the tuple $s$. In order to do this, recall the subsets $\Gamma_{q},\Sigma_{q}\subset{\mathbb{F}}^{I}$ defined in [AV22a, Sec. 6.8] and select $\gamma\in\Gamma_{q}$ and $\sigma\in\Sigma_{q}$. [^10] See also Section 5.4.
Definition 5.2.1 ([Let99, Kol14, AV22a]).
@@
$B_{i}\coloneqq F_{i}+\phi_{q}(F_{i})+\sigma_{i}k_{i}^{-1}\qquad i\in I\backslash X,$ (5.6)
@@
As a standalone algebra, $U_{q}(\mathfrak{k})$ is also called an $\imath$-quantum group, see, e.g., [BW18a, BW21, LW21, WZh23, Wa24]. $\triangledown$
@@
We will now describe a family of examples of QSPs.
@@
${\mathfrak{k}}$ =$={\mathbb{C}}\langle\{f_{i}-y_{i}e_{\tau(i)}\}_{i\in I},\,\{h_{i}-h_{\tau(i)}\}_{i\in I,\,i\neq\tau(i)}\rangle,$ (5.9)
@@ 5.3. Basic K-matrices
such that the linear relation (2.7) holds with $K=\Upsilon$ and $\psi=\phi_{q}^{-1}$. Moreover, the coproduct formula (2.9) holds with $J=R_{\phi}$, where $R_{\phi}$ denotes the universal R-matrix of the Hopf subalgebra [^11] The element $R_{\phi}$ was introduced in [AV22a, Def. 6.7, Lem. 6.8] as the R-matrix of $U_{q}({\mathfrak{g}}_{X})$ with a Cartan correction. More precisely, $R_{\phi}=(T_{\phi}^{-1}\otimes T_{\phi}^{-1})\cdot\Delta(T_{\phi})$, where $T_{\phi}$ is a modified quantum Weyl group operator for the longest element of $W_{X}$. $U_{q}({\mathfrak{g}}_{X})\cdot U_{q}({\mathfrak{h}}^{\phi})$. In particular, $(\phi_{q}^{-1},R_{\phi},\Upsilon)$ is a cylindrical structure on $(U_{q}({\mathfrak{g}}),U_{q}(\mathfrak{k}))$.
The key idea in the proof of Theorem 5.3.1 is due to H. Bao and W. Wang, who established it for certain quantum symmetric pairs in the case ${\mathfrak{g}}=\mathfrak{sl}_{N}$, see [BW18a]. M. Balagović and S. Kolb generalized it in [BK19, Secs. 6 & 9] to all quantum symmetric pairs with $\dim({\mathfrak{g}})<\infty$, obtaining a cylindrical structure with $\psi$ equal to a diagram automorphism.121212 In fact, [BK19] came close to proving Theorem 5.3.1 as stated, but, as in [BW18a], the QSP parameters were further constrained by requiring an analogue of the bar involution on $U_{q}(\mathfrak{k})$ (as observed in [Kol22], the existence of this local bar involution follows by first showing the existence of $\Upsilon$ as in [AV22a, Sec. 7] and then imposing the constraints). Also, a crucial support property of $R\cdot R_{\phi}^{-1}$ was only shown for $\dim({\mathfrak{g}})<\infty$, see [AV22a, Prop. 4.3]. Finally, to obtain the precise statement of Theorem 5.3.1, one directly checks that $\phi_{q}^{-1}$ is trivial on $U_{q}({\mathfrak{g}}_{X})\cdot U_{q}({\mathfrak{h}}^{\phi})$ and acts on the generators $B_{i}$ as a composition of the usual bar involution on $U_{q}({\mathfrak{g}})$ and an involution of the QSP parameters, see [AV22a, Prop. 8.3]. In this case, the resulting universal K-matrix acts on ${\operatorname{Mod}}_{\sf fd}(U_{q}({\mathfrak{g}}))$, yielding matrix solutions of a constant reflection equation. In [RV20] it was pointed out that the approach of [BK19] directly extends to quantum pseudo-symmetric pairs of finite type.
@@
The proof is a coideal generalization of Lusztig’s approach to the existence of the universal R-matrix, see [Lus94]. It relies on the use of certain linear maps on $U_{q}({\mathfrak{n}}^{+})$, called skew derivations, which interact nicely with the bialgebra structure of $U_{q}({\mathfrak{g}})$ as well as the pairing between $U_{q}({\mathfrak{n}}^{+})$ and $U_{q}({\mathfrak{n}}^{-})$, yielding a recursive proof of the existence of each $\Upsilon_{\lambda}$. Note that an important step in this proof, the so-called fundamental lemma of quantum symmetric pairs, first formulated in [BK15, Conj. 2.7], was uniformly established in [BW21] in the Kac-Moody setting.
@@
In [AV22a], by gauge transformations in terms of an invertible $g\in\operatorname{End}({\mathsf{f}})$ such that $\operatorname{Ad}(g)$ preserves $U_{q}({\mathfrak{g}})$, we obtain a family of related cylindrical structures. A distinguished example is given by the semistandard twist pair, which consists of the coalgebra anti-automorphism $\omega\circ\tau$ and the trivial Drinfeld twist $1\otimes 1$. It can be obtained by gauge-transforming by quantum Weyl group operators, see [AV22a, Sec. 8.10]. Note that it was essentially already described in [BK19, Sec. 7.2] with different conventions.
@@ 5.4. Generalized parameters
$\Gamma\Sigma_{q}$ =$=\big{\{}(\gamma,\sigma)\in({\mathbb{F}}^{\times})^{I}\times{\mathbb{F}}^{I}\,\big{|}\,\forall i\in I\;(\alpha_{i},\theta(\alpha_{i}))=0\;\Rightarrow\;\gamma_{i}=\gamma_{\tau(i)}\quad\text{and}$
@@ 5.5. Explicit formulae for $\Upsilon$
The construction of the element $\Upsilon$ in Theorem 5.3.1 is not explicit. If ${\mathfrak{g}}=\mathfrak{sl}_{2}$ and $\phi=\omega$, so that $U_{q}(\mathfrak{k})\subset U_{q}(\mathfrak{sl}_{2})$ deforms $U(\mathfrak{so}_{2})\subset U(\mathfrak{sl}_{2})$, one can find $\Upsilon$ by ad-hoc computations, which first appeared in [KSS93] and [tDHO98]. More generally, explicit computations have been carried out in [DK19] for all cases of restricted rank one [^14] We refer to the dimension of $({\mathfrak{h}}^{\prime})^{-\phi}$ as the restricted rank of $\phi$. in finite types with $\phi^{2}={\operatorname{id}}$.
An important consequence of the proof of Theorem 5.3.1 is that $\Upsilon_{\lambda}=0$ if $\phi(\lambda)\neq-\lambda$. In other words, $\Upsilon$ is supported only on (the nonnegative part of) the restricted root lattice, which is given by
@@
Consider the restricted Weyl group
$W^{\phi}=\{w\in W\,|\,w\circ\phi=\phi\circ w\},m$ (5.13)
which is the Weyl group of the restricted root lattice and has a presentation as a Coxeter group, see, e.g., [RV22, Sec. 4] and references therein. In [DK19], $\Upsilon$ was obtained as an ordered product of explicit K-matrices of restricted rank one. Combined with a result from [WZh23], this yields a factorization of $\Upsilon$ for all QSPs of finite type, which corresponds to a reduced expression of the longest element of $W^{\phi}$, in analogy with the factorization of the quasi-R-matrix, see [KR90, LS90].
In contrast, there is no explicit expression of universal K-matrices for quantum affine algebras. Note that, in the affine case, the resticted root system always contains all imaginary roots. In restricted rank one, the K-matrix is therefore abelian, i.e., supported only on imaginary root vectors. This observation motivates the following conjecture.
@@
In affine type, $\Upsilon$ admits an infinite-product factorization analogous to such a factorization of the (quasi-)R-matrix of quantum affine algebras and Yangians, see [KhT92].

### v2

@@ 5. Quantum symmetric pairs and K-matrices
We provide a brief summary of the theory of quantum symmetric pairs and their K-matrices, and survey some recent works on related topics. The main references for proofs and further details in this section are [Kol14, RV22, AV22, AV25b].
In [AV22, AV25b], we constructed a family of (basic and tensor) K-matrices arising from quantum symmetric pairs, whose action is defined on ${\mathcal{O}}^{\sf int}$} and, in fact, on modules in $\mathcal{O}$} satisfying a weaker integrability condition. Henceforth, for simplicity of the exposition, we consider only the category ${\mathcal{O}}^{\sf int}$}, and we set
@@ 5.1. Pseudo-involutions and their q-deformations
A Lie algebra automorphism $\theta:{\mathfrak{g}}\to{\mathfrak{g}}$ is said to be of the second kind if $\theta({\mathfrak{b}}^{+})$ is obtained by the adjoint action of the Kac-Moody group on the opposite Borel subalgebra $\mathfrak{b}$$^{-}$; this is equivalent to $\dim({\mathfrak{n}}^{+}\cap\theta({\mathfrak{n}}^{+}))<\infty$, see, e.g., [KW92, 4.6]. If $\theta$ is an involution of the second kind then, up to conjugation by an inner automorphism, we may assume $\theta({\mathfrak{h}})={\mathfrak{h}}$ and $\theta$ fixes pointwise any $\theta$-stable root space, see [KW92, Ch. 5]. The main example is the Chevalley involution $\omega:{\mathfrak{g}}\to{\mathfrak{g}}$, defined by
@@
where $\tau$ is a diagram automorphism [^9] Recall that a diagram automorphism (a bijection $\tau:I\to I$ such that $a_{\tau(i)\tau(j)}=a_{ij}$ for all $i,j\in I$) acts as a Lie algebra automorphism of $\mathfrak{g}$$^{\prime}$ by relabelling: $\tau(e_{i})=e_{\tau(i)}$, $\tau(f_{i})=f_{\tau(i)}$. By [KW92, 4.19], there exists a complementary subspace $\mathfrak{h}$$^{\prime\prime}$ of ${\mathfrak{h}}^{\prime}={\mathfrak{h}}\cap{\mathfrak{g}}^{\prime}$ such that the action of $\tau$ can be extended to ${\mathfrak{g}}={\mathfrak{g}}^{\prime}\oplus{\mathfrak{h}}^{\prime\prime}$., $\operatorname{Ad}(\widetilde{w}_{X})$ is the triple exponential action induced by the longest element $w_{X}$ of a unique subdiagram $X\subseteq I$ of finite type and a character $\chi:\mathsf{Q}\to{\mathbb{C}}^{\times}$ such that $\chi(\alpha_{i})=1$ if $i\in X$. Note that $(X,\tau,\chi)$ are uniquely determined by $\phi$ from the expression (5.3). In particular, $X=\{i\in I\,|\,\phi(\alpha_{i})=\alpha_{i}\}$ (see [RV22, Sec. 2] and [AV22, Sec. 6.1 & 6.2] for details). The case $X=\emptyset$ is referred to as quasi-split.
In order for $\phi$ to lift to an algebra automorphism $\phi_{q}:U_{q}({\mathfrak{g}})\to U_{q}({\mathfrak{g}})$, it suffices that the extension of $\tau$ to $\mathfrak{h}$$^{\prime\prime}$ described by [KW92, 4.19] permutes the $\mathbb{Z}$}-basis of $\mathsf{Q}^{\vee}_{\sf ext}$, constraining the choice of $d_{s}$ ($1\leq s\leq{\operatorname{cork}}(A)$). In the affine case, a suitable choice of $\mathbb{Z}$}-basis can always be made, see [Kol14, Prop. 2.12], although not in general if ${\operatorname{cork}}(A)=1$, see [AV22, Sec. 3.15].
In [AV22, Sec. 6.7] a definition for the lift $\phi_{q}:U_{q}({\mathfrak{g}})\to U_{q}({\mathfrak{g}})$ is given, which can be related to the original constructions in [Let99, Kol14]. Note that here we include into $\phi_{q}$ the adjoint action of a lift of $\chi$ to a multiplicative character $\chi_{q}:\mathsf{Q}\to{\mathbb{F}}^{\times}$ such that $\chi_{q}(\alpha_{i})=1$ if $i\in X$. In terms of the formalism of unrestricted specialization, see, e.g., [DCK90], the automorphism $\phi_{q}$ specializes to the automorphism $\phi$.
@@ 5.2. Pseudo-fixed-point subalgebras and their q-deformations
It is useful to constrain $\phi$ by imposing additional conditions on $X(\phi)$, $\tau(\phi)$, $\chi(\phi)$, leading to the notion of an (enriched) generalized Satake diagram, see [RV22, Def. 3.4] and [AV22, Sec. 6.4]. The resulting conditions on $\phi$ are equivalent to requiring that the associated pseudo-fixed-point subalgebra
@@
It is convenient to allow some extra degrees of freedom and define a modification $U({\mathfrak{k}})_{s}$ of $U({\mathfrak{k}})$ by adding a scalar $s_{i}\in{\mathbb{C}}$ to the generator $b_{i}\in U({\mathfrak{k}})$ for $i\in I\backslash X$. In order to define the Letzter-Kolb deformation of $U({\mathfrak{k}})_{s}$, we choose a pseudo-involution $\phi$ with a canonical expression (5.3) in terms of an associated enriched generalized Satake diagram $(X,\tau,\chi)$. The induced constraint on $\chi_{q}$ is conveniently rewritten in terms the tuple $\gamma=(\chi_{q}(\alpha_{\tau(i)}))_{i\in I}$. Also, to deduce basic properties of the deformed algebra, we will need a constrained tuple $\sigma\in{\mathbb{F}}^{I}$ which specializes to the tuple $s$. In order to do this, recall the subsets $\Gamma_{q},\Sigma_{q}\subset{\mathbb{F}}^{I}$ defined in [AV22, Sec. 6.8] and select $\gamma\in\Gamma_{q}$ and $\sigma\in\Sigma_{q}$. [^10] See also Section 5.4.
Definition 5.2.1 ([Let99, Kol14, AV22]).
@@
$\X,B_{i}\coloneqq F_{i}+\phi_{q}(F_{i})+\sigma_{i}k_{i}^{-1}\qquad i\in I\backslash$ X, (5.6)
@@
As a standalone algebra, $U_{q}(\mathfrak{k})$ is also called an $\imath$-quantum group, see, e.g., [BW18b, BW21, LW21, WZh23, Wa24]. $\triangledown$
@@
We now explicitly describe a family of examples of QSPs built from a quasi-split symmetric pair.
@@
${\mathfrak{k}}$ $={\mathbb{C}}\langle\{f_{i}-y_{i}e_{\tau(i)}\}_{i\in I},\,\{h_{i}-h_{\tau(i)}\}_{i\in I,\,i\neq\tau(i)}\rangle,$ (5.9)
@@ 5.3. Basic K-matrices
    such that the linear relation (2.7) holds with $K=\Upsilon$ and $\psi=\phi_{q}^{-1}$.
  2. (2)
@@
    The coproduct formula (2.9) holds with $J=R_{\phi}$, where $R_{\phi}$ denotes the universal R-matrix of the Hopf subalgebra $U_{q}({\mathfrak{g}}_{X})\cdot U_{q}({\mathfrak{h}}^{\phi})$. [^11] The element $R_{\phi}$ was introduced in [AV22, Def. 6.7, Lem. 6.8] as the R-matrix of $U_{q}({\mathfrak{g}}_{X})$ with a Cartan correction. More precisely, $R_{\phi}=(T_{\phi}^{-1}\otimes T_{\phi}^{-1})\cdot\Delta(T_{\phi})$, where $T_{\phi}$ is a modified quantum Weyl group operator for the longest element of $W_{X}$.
In particular, $(\phi_{q}^{-1},R_{\phi},\Upsilon)$ is a cylindrical structure on $(U_{q}({\mathfrak{g}}),U_{q}(\mathfrak{k}))$, called standard.
As stated above, the theorem was proved in [AV22], building on the following earlier results. H. Bao and W. Wang established (1) in [BW18a, Ch. 2] for the family of quasi-split quantum symmetric pairs with ${\mathfrak{g}}=\mathfrak{sl}_{N}$ and $\tau$ nontrivial. M. Balagović and S. Kolb generalized (1) to all quantum symmetric pairs in [BK19, Sec. 6]. [^12] In [BW18a, BK19], the QSP parameters were constrained by requiring an analogue of the bar involution on $U_{q}(\mathfrak{k})$, which was directly involved in the defining equation $\Upsilon$. In our approach, (2.7) takes the form $\Upsilon\cdot b=\phi_{q}^{-1}(b)\cdot\Upsilon$ for any $b\in U_{q}(\mathfrak{k})$. Then, a linear relation generalizing the one from [BW18a, BK19] appears by observing that $\phi_{q}^{-1}$ is trivial on $U_{q}({\mathfrak{g}}_{X})\cdot U_{q}({\mathfrak{h}}^{\phi})$ and acts on the generators $B_{i}$ as a composition of the usual bar involution on $U_{q}({\mathfrak{g}})$ and an involution of the QSP parameters, see [AV22, Prop. 8.3]. Based on this observation, the existence of a QSP bar involution was later obtained in [Kol22] from the existence of $\Upsilon$, as defined by (2.7), and then imposing the constraints. The proof of (2) appeared in [BK19, Sec. 9] for finite types. The general case was proved in [AV22] and relied on a crucial support property of $R\cdot R_{\phi}^{-1}$, see [AV22, Prop. 4.3].
Finally, in [RV20] it was pointed out that the approach of [BW18a, BK19] directly extends to quantum pseudo-symmetric pairs of finite type and the results of [AV22] are proved in this generality.
@@
The proof is a generalization of Lusztig’s approach to the universal R-matrix, see [Lus94]. The key tool is given by skew derivations, certain linear maps on $U_{q}({\mathfrak{n}}^{+})$ which interact nicely with the bialgebra structure of $U_{q}({\mathfrak{g}})$, as well as the pairing between $U_{q}({\mathfrak{n}}^{+})$ and $U_{q}({\mathfrak{n}}^{-})$, yielding a recursive proof of the existence of each $\Upsilon_{\lambda}$. An important step, the so-called fundamental lemma of quantum symmetric pairs, first formulated in [BK15, Conj. 2.7], was uniformly established in [BW21] in the Kac-Moody setting.
@@
In [BK19], gauge transformations are implicitly used to obtain a cylindrical structure for any QSP with $\dim({\mathfrak{g}})<\infty$ and $\psi$ equal to a diagram automorphism. In this case, the resulting universal K-matrix acts on ${\operatorname{Mod}}_{\sf fd}(U_{q}({\mathfrak{g}}))$, yielding matrix solutions of a constant reflection equation. By gauge transformations in terms of an invertible $g\in\operatorname{End}({\mathsf{f}})$ such that $\operatorname{Ad}(g)$ preserves $U_{q}({\mathfrak{g}})$, we obtain a family of related cylindrical structures. A distinguished example is given by the semistandard twist pair, which consists of the coalgebra anti-automorphism $\omega\circ\tau$ and the trivial Drinfeld twist $1\otimes 1$. It can be obtained by gauge-transforming by quantum Weyl group operators, see [AV22, Sec. 8.10]. Note that it was essentially already described in [BK19, Sec. 7.2], up to conventions.
@@ 5.4. Generalized parameters
$\Gamma\Sigma_{q}$ $=\big\{(\gamma,\sigma)\in({\mathbb{F}}^{\times})^{I}\times{\mathbb{F}}^{I}\,\big|\,\forall i\in I\;(\alpha_{i},\theta(\alpha_{i}))=0\;\Rightarrow\;\gamma_{i}=\gamma_{\tau(i)}\quad\text{and}$
@@ 5.5. Explicit formulae for $\Upsilon$
The construction of the element $\Upsilon$ in Theorem 5.3.1 is not explicit. If ${\mathfrak{g}}=\mathfrak{sl}_{2}$ and $\phi=\omega$, so that $U_{q}(\mathfrak{k})\subset U_{q}(\mathfrak{sl}_{2})$ deforms $U(\mathfrak{so}_{2})\subset U(\mathfrak{sl}_{2})$, one can find $\Upsilon$ by ad-hoc computations, which first appeared in [KSS93] and [tDHO98]. More generally, explicit computations have been carried out in [DK19] for many cases of restricted rank one [^14] We refer to the dimension of $({\mathfrak{h}}^{\prime})^{-\phi}$ as the restricted rank of $\phi$. in finite types with $\phi^{2}={\operatorname{id}}$.
An important consequence of the proof of Theorem 5.3.1 is that $\Upsilon_{\lambda}=0$ if $\phi(\lambda)\neq-\lambda$. In other words, $\Upsilon$ is supported on the nonnegative part of the restricted root lattice. The latter is given by
@@
Consider the restricted Weyl group $\widetilde{W}$, i.e. the Weyl group of the restricted root lattice. It has a presentation as a Coxeter group, see, e.g., [RV22, Sec. 4] and references therein. In [DK19], $\Upsilon$ was obtained as an ordered product of explicit K-matrices of restricted rank one. Combined with a result from [WZh23], this yields a factorization of $\Upsilon$ for all QSPs of finite type, which corresponds to a reduced expression of the longest element of $\widetilde{W}$, in analogy with the factorization of the quasi-R-matrix in finite types, see [KR90, LS90].
In contrast, there is no explicit expression of universal K-matrices for quantum affine algebras.
@@
Given a quantum symmetric pair of affine type, $\Upsilon$ admits a loop-triangular factorization analogous to that of the (quasi-)R-matrix of quantum affine algebras, see [KhT92, LSS93, Be94, Dam98].
For affine symmetric pairs of restricted rank one, the K-matrix is abelian, i.e., supported only on imaginary root vectors. Hence, the conjecture is obviously true in this case. Note that, the resticted root system always contains all imaginary roots. Thus, the abelian factor is always expected to be non-trivial. We also expect that each factor is expressed by an infinite product of restricted rank one K-matrices, indexed by a convex order on the positive restricted root system.